# Exploring Feedback Generation in Automated Skeletal Movement Assessment: A Comprehensive Overview

### v3

@@
Abstract—The application of machine-learning solutions to movement assessment from skeleton videos has attracted significant research attention in recent years. This advancement has made rehabilitation at home more accessible, utilizing movement assessment algorithms that can operate on affordable equipment for human pose detection and analysis from 2D or 3D videos. While the primary objective of automatic assessment tasks is to score movements, the automatic generation of feedback highlighting key movement issues has the potential to significantly enhance and accelerate the rehabilitation process. While numerous research works exist in the field of automatic movement assessment, only a handful address feedback generation. In this study, we explain the types of feedback that can be generated, review existing solutions for automatic feedback generation, and discuss future research directions. To our knowledge, this is the first comprehensive review of feedback generation in skeletal movement assessment.
@@ 1 Introduction
While numerous research works exist in the field of automatic movement assessment, only a handful provide algorithms for feedback generation. This paper aims to address this gap by examining the nature and significance of feedback within the context of movement assessment. Specifically, we will explore the types of feedback that can be generated, discuss the approaches and challenges involved in feedback generation, and review existing solutions in this area.
@@ 3 Feedback Generation
Defining ”feedback” can be complex. In the context of movement assessment, it encompasses a wide array of response types, occurring during or after a performance. Feedback may consist of both basic and advanced guidance, conveyed verbally [3, 28, 33], textually [1, 24, 33, 31, 22], or visually [3, 1, 12, 47, 28, 40], and occasionally through physical means [21]. Table 1 provides a summary of feedback generation solutions, based on the criteria proposed in the following sections.
Scoring Model.
@@
The primary objective in movement assessment is score generation. Typically, scoring is accomplished through regression or classification machine-learning algorithms, trained with sample data to minimize prediction errors [23]. These algorithms often derive features from various subsets of raw input parameters or other features, aiding in prediction accuracy. However, interpreting these features can be challenging, making it less immediate to identify performance errors and generate corresponding feedback. Alternatively, models that solely utilize raw input parameters may offer clearer interpretability, enabling a more straightforward identification of performance parameters contributing most to the score, which can then be translated into feedback [24]. Furthermore, such approaches may be more adaptive in assessing and providing feedback for new movement types without prior assumptions. In cases where this is not feasible, another statistical model may be necessary to predict or detect mistakes and translate them into feedback [31, 47].
Feedback Boundness and Order.
@@
According to [23], feedback can be categorized as either bound or unbound. Feedback is considered bound when it is limited to predefined mistakes that can be detected [31, 47]. Conversely, unbound feedback may encompass detected and interpreted mistakes in any provided parameter, translating them into text or visual representation [25]. Another parameter used to classify feedback types is the feedback order. First-order feedback mainly addresses mistakes related to a single parameter at a specific point in time. Second and third-order feedback may involve temporal and skeletal aggregations, resulting in more informative output, as illustrated in Figure 1. The advantage of bound feedback lies in the ability of the prediction model to be trained directly to predict higher-order feedback [31, 47]. In contrast, with unbound feedback, the aggregation step from first-order feedback to higher-order feedback is not always straightforward [24]. Higher-order feedback may include guidelines for addressing mistakes, which may not necessarily be linked to the raw parameters representing the mistakes directly, thus requiring a system with access to prior knowledge. Recent advancements in Natural Language Processing (NLP), such as Large Language Models (LLMs) [51, 13], including tools like ChatGPT, may facilitate the development of new types of feedback.
Figure 1: Aggregating first-order feedback. Each cell painted red with a sign represents a first-order deviation or first-order feedback. However, if a parameter deviates consecutively or nearly consecutively, temporal aggregation can introduce feedback of a higher order. Similarly, when parameters deviate simultaneously, skeletal aggregation can introduce feedback of an even higher order.
Temporal Granularity.
@@
Figure 2: Different approaches to adding temporal notation to bound feedback labels. In (a) different feedback labels are detected along the movement time axis using sliding windows of different lengths. In (b) a set of time frames is defined to support temporally local feedback generation. In (c) Temporally local feedback labels are achieved using a Cartesian product of the sets of feedback labels and time frames.
Offline/Online Feedback Generation Mode.
@@
Feedback Modality.
Creating textual or verbal feedback that is both intuitive and informative presents challenges, as discussed earlier. In contrast, visual feedback holds the potential to be more intuitive and informative, and sometimes even easier to generate. For instance, presenting multiple first-order unbound feedback labels simultaneously as an animation on a skeleton figure inherently incorporates skeletal and temporal aggregation, as illustrated in Figure 3. As previously discussed, several works have proposed solutions for visual feedback generation.
Figure 3: A proposed method for offering highly relevant and intuitive feedback by visualizing only the first-order deviations together. In this example, although the three reported deviations are first-order, which could be overwhelming in textual form, the visualization clearly indicates that the main issue is the right elbow being too low. In other words, skeletal aggregation occurs naturally through perception.
@@ 4 Evaluating Generated Feedback
In predicting unbound feedback, the aggregation algorithms that elevate the feedback order may significantly influence the quality evaluation of the generated feedback. Similarly, heuristics and temporal aggregation algorithms that transform temporally local feedback labels into global ones may also play a crucial role in this evaluation.
Bound Feedback Labels.
@@
Unbound Feedback Labels.
Similar to generated unbound feedback, the ground-truth labels for unbound feedback may require inclusion of the set of deviated parameters, their corresponding types of deviations, and a temporal bounding box. When ground-truth feedback labels lack skeletal aggregation, each parameter and its deviation are denoted as separate feedback labels. Conversely, when feedback labels include skeletal aggregation, all involved parameters are encompassed within a single label. Unlike bound feedback labels, annotating unbound feedback labels can be an exhaustive task, as each parameter must be individually assessed by a professional, considering aggregations as well. However, these annotations may solely be utilized for feedback evaluation, as unbound feedback is typically generated from aggregated first-order deviations, which do not necessarily require training on feedback labels.
Skeletal IoU.
Similar to how temporal IoU is employed as a metric for temporal localization, skeletal IoU should be utilized to assess the quality of generated unbound feedback. This involves evaluating the similarity between the predicted and ground-truth sets of involved parameters and their deviations. Specifically, this entails calculating a discrete IoU between two sets of deviated parameters, denoted as $P$ and $G$, representing the predicted and ground-truth deviated parameters, respectively. Each deviated parameter consists of a tuple indicating the parameter identity and the deviation type. Essentially, this calculation represents the Jaccard similarity between the two sets,
@@
Online-Feedback Evaluation.
@@ 5 Discussion
In this review, we aimed to explore the definition of feedback within the context of automatic movement assessments, along with its various forms and the challenges associated with providing solutions for each. We have addressed several aspects in which solutions may vary, with particular emphasis on the concept of feedback boundness.
@@
In upcoming studies focusing on feedback generation, adopting the terminology outlined in this paper can be beneficial. For instance, in a proposed solution generating bound feedback, it is beneficial to mention when the feedback labels are of high order and if a quantitative evaluation is included. Furthermore, acknowledging the need for numerous training samples for each feedback label can justify a relatively limited number of supported feedback labels. Similarly, in a proposed solution generating unbound feedback, it is valuable to underscore feedback versatility while mentioning the impracticality of annotating feedback labels, which could justify an inclusion of qualitative evaluation only.
@@ 6 Conclusions
In this work, we have emphasized the significance of feedback generation in automatic movement assessment solutions. We have explored various types of feedback that can be generated and discussed the associated challenges, including data annotation, algorithm development, and evaluation. Our aim is for the insights from this work to inform better design decisions in future systems, facilitate accessibility to feedback generation, and contribute to standardizing the terminology used in the literature.
@@ Acknowledgement
I dedicate this work to our beloved dog, Benji, who sadly passed away on April 10, 2024, at the age of 16 years and 8 months, while this study was in progress. Benji had a pure soul and was a true friend, playing a significant role in our lives, particularly in mine. He was always by my side, providing companionship and support throughout much of my career journey. Unfortunately, during his final month, he battled heart and liver disease along with other health issues, compromising his quality of life. Despite our efforts to aid his recovery, it became evident that his suffering was too great, and we made the difficult decision to euthanize him. My dear Benji, losing you after all these years together has been so traumatic. The realization that you are no longer around and that I will never see you again is heartbreaking. I love you and I will never forget you. May your soul rest in peace.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]

### v4

# Evaluation Framework for Feedback Generation Methods in Skeletal Movement Assessment

@@
Abstract
The application of machine-learning solutions to movement assessment from skeleton videos has attracted significant research attention in recent years. This advancement has made rehabilitation at home more accessible, utilizing movement assessment algorithms that can operate on affordable equipment for human pose detection and analysis from 2D or 3D videos. While the primary objective of automatic assessment tasks is to score movements, the automatic generation of feedback highlighting key movement issues has the potential to significantly enhance and accelerate the rehabilitation process. While numerous research works exist in the field of automatic movement assessment, only a handful address feedback generation. In this study, we propose terminology and criteria for the classification, evaluation, and comparison of feedback generation solutions. We discuss the challenges associated with each feedback generation approach and use our proposed criteria to classify existing solutions. To our knowledge, this is the first work that formulates feedback generation in skeletal movement assessment.
@@ 1 Introduction
While numerous research works exist in the field of automatic movement assessment, only a handful provide algorithms for feedback generation. This paper aims to address this gap by examining the nature and significance of feedback within the context of movement assessment. Specifically, we will explore the types of feedback that can be generated, discuss the approaches and challenges involved in feedback generation, and use the proposed terminology to classify existing solutions in this area.
This work addresses feedback generation as a computational problem and does not cover the deployment process in clinical scenarios.
@@ 3 Feedback Generation
Defining “feedback” can be complex. In the context of movement assessment, it encompasses a wide array of response types, occurring during or after a performance. Feedback may consist of both basic and advanced guidance, conveyed verbally [3, 28, 33], textually [1, 24, 33, 31, 22], or visually [3, 1, 12, 47, 28, 40], and occasionally through physical means [21]. Table 1 provides a summary of feedback generation solutions, based on the criteria proposed in the following sections.
3.1 Scoring Model
@@
The primary objective in movement assessment is score generation. Typically, scoring is accomplished through regression or classification machine-learning algorithms, trained with sample data to minimize prediction errors [23]. These algorithms often derive compound features from various subsets of raw geometric pose and movement features, or handcrafted features, aiding in prediction accuracy. However, interpreting compound features can be challenging, making it less immediate to identify performance errors and generate corresponding feedback. Alternatively, models that solely utilize raw features may offer clearer interpretability, enabling a more straightforward identification of feature deviations that have the highest impact on the score, which can then be translated into feedback [24]. For instance, a single component of a single 3D joint location vector, or velocity vector, may be a raw feature. For example, a deviation of the Y component of one of the elbows can be translated directly to feedback regarding the elbow’s height. The raw features approach may be more adaptive in assessing and providing feedback for new movement types without prior assumptions, due to their general nature. In cases where using raw features is not feasible, another statistical model may be necessary to predict or detect mistakes and translate them into feedback [31, 47].
3.2 Feedback Boundness and Order
@@
According to [23], feedback can be categorized as either bound or unbound. Feedback is considered bound when it is limited to predefined mistakes that can be detected [31, 47]. Conversely, unbound feedback may encompass detected and interpreted mistakes, or deviations, in any provided raw feature, translating them into text or visual representation [25]. Another parameter used to classify feedback types is the feedback order. First-order feedback labels refer to mistakes, or deviations, related to a single raw feature at a specific point in time. Second and third-order feedback labels may involve temporal and skeletal aggregations, resulting in more abstract and intuitive output. When only one of the two aggregation types is applied, the feedback labels will be considered second-order, whereas when both are applied, they will be considered third-order. Raw feature deviations are demonstrated in Figure 1. The advantage of bound feedback lies in the ability of the prediction model to be trained directly to predict higher-order feedback [31, 47]. This is possible by addressing the bound feedback generation problem as a detection or classification problem, where the set of predefined labels can include abstract, high-order labels. Such labels, for example, can include information regarding balance, strength, number of repetitions, or other details that may not be easily derived from raw feature deviations. Meanwhile, with unbound feedback, the aggregation step from first-order feedback to higher-order feedback is not always straightforward and may require the design of a problem formulation, which may be addressed using heuristics or optimization problems [24]. Recent advancements in Natural Language Processing (NLP), such as Large Language Models (LLMs) [51, 13], including tools like ChatGPT, might allow an automated feedback aggregation, which will not be constrained by any proposed formulation.
Figure 1: Examples of raw feature deviations. A green, clear cell indicates a proper, in-range feature value, while cells highlighted in red indicate deviating feature values. Each deviation can be directly translated into a first-order feedback label. When a feature consistently deviates over a sequence of frames, temporal aggregation will consolidate the sequence of first-order feedback labels into a single feedback label. Similarly, when related features deviate simultaneously, skeletal aggregation will consolidate the set of related first-order feedback labels into a single feedback label.
3.3 Temporal Granularity
@@
Figure 2: Different approaches to adding temporal notation to bound feedback labels: (a) different feedback labels are detected along the movement time axis using sliding windows of different lengths; (b) a set of time frames is defined to support temporally local feedback generation; and (c) temporally local feedback labels are achieved using a Cartesian product of the sets of feedback labels and time frames.
3.4 Offline/Online Feedback Generation Mode
@@
3.5 Feedback Modality
Creating textual feedback that is both intuitive and informative presents challenges, as discussed earlier. In most cases, verbal feedback will be derived from textual feedback using text-to-speech technologies, and thus it encounters similar challenges. In contrast, visual feedback holds the potential to be more intuitive and informative, and sometimes even easier to generate. For instance, presenting multiple first-order unbound feedback labels simultaneously as an animation on a skeleton figure inherently incorporates skeletal and temporal aggregation, as illustrated in Figure 3.
Figure 3: Visualizing multiple first-order feedback labels in (a), versus a single second-order feedback label in (b). This example demonstrates the power of visual feedback over textual feedback: the three reported deviations in (a) are first-order, yet, they provide intuitive feedback, leaving negligible added value for the aggregated feedback label in (b). This approach has the potential to enable unbound feedback generation without addressing the skeletal aggregation problem.
@@ 4 Evaluating Generated Feedback
When predicting unbound feedback, the aggregation algorithms that elevate the feedback order may significantly influence the quality evaluation of the generated feedback. Similarly, heuristics and temporal aggregation algorithms that transform temporally local feedback labels into global ones may also play a crucial role in this evaluation.
4.1 Bound Feedback Labels
@@
4.2 Unbound Feedback Labels
Similar to generated unbound feedback, the ground-truth labels for unbound feedback may require inclusion of the set of deviated features, their corresponding types of deviations, and a temporal bounding box. When ground-truth feedback labels lack skeletal aggregation, each feature and its deviation are denoted as separate feedback labels. Conversely, when feedback labels include skeletal aggregation, all involved features are encompassed within a single label. Unlike bound feedback labels, annotating unbound feedback labels can be an exhaustive task, as each feature must be individually assessed by a professional, considering aggregations as well. However, these annotations may solely be utilized for feedback evaluation, as unbound feedback is typically generated from aggregated first-order deviations, which do not necessarily require training on feedback labels.
4.3 Feature IoU
In the case of unbound feedback generation, temporal IoU is a common means for evaluating the temporal accuracy of temporally-aggregated feedback labels. Likewise, when applying skeletal aggregation, feature IoU should be utilized to measure the quality of generated feedback labels. This involves evaluating the similarity between the predicted and ground-truth sets of involved features and their deviations. Specifically, this entails calculating a discrete IoU between two sets of deviated features, denoted as $P$ and $G$, representing the predicted and ground-truth deviated features, respectively. Each deviated feature consists of a tuple indicating the feature identity and the deviation type. Essentially, this calculation represents the Jaccard similarity between the two sets,
@@ 4.3 Feature IoU
4.4 Online-Feedback Evaluation
@@ 5 Discussion
In this study, we aimed to explore the definition of feedback within the context of automatic movement assessment, along with its various forms and the challenges associated with providing solutions for each. We have addressed several aspects in which solutions may vary, with particular emphasis on the concept of feedback boundness.
@@
In upcoming studies focusing on feedback generation, adopting the terminology and criteria outlined in this paper can be beneficial. For instance, in a proposed solution generating bound feedback, it is beneficial to mention when the feedback labels are of high order and if a quantitative evaluation is included. Furthermore, acknowledging the need for numerous training samples for each feedback label might justify a relatively limited number of supported feedback labels. Similarly, in a proposed solution generating unbound feedback, it is valuable to underscore feedback versatility while mentioning the impracticality of annotating feedback labels, which might justify an inclusion of qualitative evaluation only.
@@ 6 Conclusions
This study addresses the critical gap in automatic movement assessment by focusing on the generation of feedback, which is essential for enhancing and accelerating the rehabilitation process. By proposing a clear terminology and criteria for the classification, evaluation, and comparison of feedback generation solutions, we provide a structured approach to tackling the challenges inherent in this area. Our classification framework helps in understanding and improving existing solutions, promoting the development of more effective and interpretable feedback mechanisms. As the first work to formulate feedback generation in skeletal movement assessment, our research lays the groundwork for future advancements, aiming to make rehabilitation more accessible and efficient through sophisticated machine-learning algorithms.
@@ Acknowledgement
I dedicate this work to our beloved dog, Benji, who sadly passed away while this study was in progress, at the age of 16 years and 8 months. Benji had a pure soul and was a true friend, playing a significant role in our lives, particularly in mine. He was always by my side, providing companionship and support throughout much of my career journey. Unfortunately, during his final month, he battled heart and liver disease along with other health issues, compromising his quality of life. Despite our efforts to aid his recovery, it became evident that his suffering was too great, and we made the difficult decision to euthanize him. My dear Benji, losing you after all these years together has been so traumatic. The realization that you are no longer around and that I will never see you again is heartbreaking. I love you and I will never forget you. May your soul rest in peace.